# Towards Quantum Computing Timelike Hadronic Vacuum Polarization and Light-by-Light Scattering: Schwinger Model Tests

### v1

@@ I Introduction
State of the art theoretical calculations of the QCD contributions involve either the use of dispersive methods or euclidean lattice approaches [4]. Dispersive methods make use of the fact that a virtual photon decaying to hadrons can be related to $e^{+}e^{-}$ annihilation to QCD states, and that the hadronic spectral functions can be extracted from QCD decays. As a result, this is an intrinsically data driven approach. In contrast, in lattice based calculations one can directly extract correlators of electromagnetic currents, which are related to the HVP and HLBL. Although lattice QCD is the only true first-principle approach available, it is limited to computations of correlators of space-like separated currents.
In this work we propose using quantum computers (QCs) to circumvent some of the issues found in Euclidean lattice calculations. The main motivation to using QCs is that they allow for direct computations in Minkowski time without the presence of a sign problem for any spacetime separation. In the current stage of QC hardware, however, the available quantum devices are still too small and error-prone to effectively perform any meaningful calculations in QCD, and any precision calculation, if possible, will only be achieved in some far future [5, 6, 7]. As a result, here we illustrate the approach in 1+1-d QED, i.e. the Schwinger model [8]. This theory shares several phenomenological properties with 3+1-d QCD (see e.g. [9, 10, 11]), but it is technically simpler to implement on the lattice [12, 13]. In particular, as we detail below, it can be efficiently analyzed using (classical) tensor network methods, see e.g. [14], as a precursor to implementations in large scale digital QCs [15, 16]. We recall that, despite the classical nature of tensor networks, their implementation mimics several aspects faced when using QCs, and thus both approaches are intimately connected, as we illustrate below.
@@ III Hadronic vacuum polarization
In Fig. 3 we show the real-time quantum simulation of $\Pi^{00}$ (in position space) for $\Delta x=x-x_{0}=2$ in lattice units, in a smaller system with $N=20$ qubits.
@@ Acknowledgements.
We thank A. Florio, R. Pisarski, A. V. Sadofyev, and R. Szafron for helpful discussions related to this work. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics and National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704. The work of J.R. was supported by the U.S. Department of Energy, Office of Science, Office of Workforce Development for Teachers and Scientists (WDTS) under the Science Undergraduate Laboratory Internships Program (SULI) and the Brookhaven National Laboratory (BNL), Physics Dept. under the BNL Supplemental Undergraduate Research Program (SURP).

### v2

# Towards a Real-Time Computation of Timelike Hadronic Vacuum Polarization and Light-by-Light Scattering: Schwinger Model Tests

@@ I Introduction
State of the art theoretical calculations of the QCD contributions involve either the use of dispersive methods or euclidean lattice approaches [4]. Dispersive methods make use of the fact that a virtual photon decaying to hadrons can be related to $e^{+}e^{-}$ annihilation to QCD states, and that the hadronic spectral functions can be extracted from QCD decays. As a result, this is an intrinsically data driven approach. In contrast, in lattice based calculations one can directly extract correlators of electromagnetic currents, which are related to the HVP and HLBL. Traditional lattice QCD calculations are carried out in Euclidean time and, therefore, only limited HVP with spacelike momenta transfers. The spacelike lattice QCD results can be analytically continued to the timelike region only within the kinematic domains where no physical intermediate states are allowed. For the HVP this domain is limited by twice the pion mass, which vanishes in the chiral limit, for more detailed discussions see [5, 6]. On the other hand, presently there are significant discrepancy between the results of hadronic contribution to HVP from lattice QCD calculations [7] and those estimated from the fits to the multiple experimental results over an extended timelike region [1]. Resolving this discrepancy is critical in determining whether BSM physics contribute to the muon’s $g-2$. In light of this, it is very much welcome to have non-perturbative theoretical technique that can not only access HVP in an extended timelike region and at the same time capable of continuing to the spacelike region.
In this work we propose using quantum computers (QCs) to circumvent some of the issues found in Euclidean lattice calculations. The main motivation to using QCs is that they allow for direct computations in Minkowski time without the presence of a sign problem for any spacetime separation. In the current stage of QC hardware, however, the available quantum devices are still too small and error-prone to effectively perform any meaningful calculations in QCD, and any precision calculation, if possible, will only be achieved in some far future [8, 9, 10]. As a result, here we illustrate the approach in 1+1-d QED, i.e. the Schwinger model [11]. This theory shares several phenomenological properties with 3+1-d QCD (see e.g. [12, 13, 14]), but it is technically simpler to implement on the lattice [15, 16]. In particular, as we detail below, it can be efficiently analyzed using (classical) tensor network methods, see e.g. [17], as a precursor to implementations in large scale digital QCs [18, 19]. [^1] See for example [20, 21, 22, 23, 24, 25, 26] for further related studies of the gauge theories using quantum computing and tensor ntwork methods. We recall that, despite the classical nature of tensor networks, their implementation mimics several aspects faced when using QCs, and thus both approaches are intimately connected, as we illustrate below.
@@ III Hadronic vacuum polarization
In Fig. 3 we show the real-time quantum simulation of $\Pi^{00}$ (in position space) for $\Delta x=x-x_{0}=2$ in lattice units, in a smaller system with $N=20$ qubits. Unlike the above classical real time simulations, where the ground state can be directly obtained using the DMRG algorithm, constructing the ground for a quantum simulation is, in general, at least as hard as performing the subsequent real time evolution. In the present simulations we circumvent this issue by performing the simulation using the built-in classical emulator of a quantum computer in Qiskit. This allows the ground state to be directly converted into a quantum circuit, while the state itself can be directly obtained for the small system we consider. [^4] By this we mean that in a small system, the current approach is equivalent to directly diagonalizing the Hamiltonian, and thus we have access to the complete Hilbert space. Alternatively, we could have, for this small system, loaded the state from a DMRG run using tensor networks, or more generally used a hybrid variational method, such as VQE, to obtain this state.
@@
We again recall that the transformation to momentum space can be done the same way as shown for the above MPS example. However, one important difference compared to tensor network methods relates to the extraction of the expectation value of hermitian operators, since quantum computers only naturally allow for the application of unitary gates (operators). In the example shown in Fig. 3 this is done by using the simple interference measurement protocol, i.e. the Hadamard test, illustrated in Fig. 4. Taking the initial state of a quantum simulation, $|\Omega\rangle$, and adding a single ancilla qubit in the $|+\rangle=(|0\rangle+|1\rangle)/\sqrt{2}$, one can apply a series of gates to generate a characteristic interfering pattern:
@@
where $U=e^{-itH}$ is the time evolution operator and $Z_{m}$ denotes the relevant operator in the lattice discretiz¡tion of the current operators. The evolution in Eq. (6) is implemented by the first four gates in Fig. 4. The last two gates in the circuit diagram implement the measurement of the ancilla in the $X$-basis, which allow to directly compute $\text{Re}\bra{\Omega}Z_{m}(t)Z_{n}\ket{\Omega}$. A Qiskit code to compute this function is provided in [40], along with more details.
@@ Acknowledgements.
We thank A. Florio, R. Pisarski, A. V. Sadofyev, and R. Szafron for helpful discussions related to this work. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics and National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704. The work of S.M. is partially supported by Laboratory Directed Research and Development (LDRD) funds from Brookhaven Science Associates. The work of J.R. was supported by the U.S. Department of Energy, Office of Science, Office of Workforce Development for Teachers and Scientists (WDTS) under the Science Undergraduate Laboratory Internships Program (SULI) and the Brookhaven National Laboratory (BNL), Physics Dept. under the BNL Supplemental Undergraduate Research Program (SURP).